OpenStudy (sesca89):

The human body has a volume of 2 liters of red blood cells. The volume of one red blood cell is 9 x 10^-14 liters. Approximately how many red blood cells are in the human body? Express your answer in scientific notation.

OpenStudy (sesca89):

\[9 \times 10 ^{-14}\] that's what it's supposed to look like

OpenStudy (perl):

did you try division

OpenStudy (sesca89):

Well...yes..but i got ≈ 2.2222222222 • 10^14 and that was wrong

OpenStudy (sesca89):

hold on

OpenStudy (sesca89):

\[\approx 2.222222222 \times 10^{14}\]

OpenStudy (perl):

should be 2.22 x 10^13

OpenStudy (sesca89):

okay, can you tell me how you found that?

OpenStudy (sesca89):

Oh, well thanks!

OpenStudy (mrnood):

if 1 cell is 9*10^-14 l then 1 litre is 1/9*10^-14 = 1/9 *10^14 = 0.11 *10 ^14 so 2l = 0.22 *10^14 = 2.2 *10^13